# WOD-E2E: Waymo Open Dataset for End-to-End Driving in Challenging Long-tail Scenarios

### v2

@@ 4.3 Benchmark Models
Swin-Trajectory [24] is a lightweight, MLP-based model. It uses a Swin Transformer [18] to extract image features from three front cameras and a simple MLP to directly predict waypoints. The model is lightweight and achieves a slightly better RFS (7.543) than the baseline.
DiffusionLTF and UniPlan are both Diffusion-based models built on the DiffusionDrive [17] architecture. Their primary difference lies in the training datasets used: DiffusionLTF utilizes WOD-E2E, CARLA [9], NAVSIM [8], and WOD-Perception [26], whereas UniPlan is trained on WOD-E2E and nuPlan [3]. They achieve comparable performance, with RFS scores of 7.717 and 7.779, respectively.
Poutine [21], HMVLM [29], and AutoVLA [40] are all MLLM-based models that use Qwen2.5 as their backbone. They share a similar problem formulation, taking camera images and ego states as input modalities and outputting future waypoints as text. Additionally, all three models utilize Chain-of-Thought (CoT) reasoning before generating a trajectory. Despite these similarities, their results show a significant performance gap, with AutoVLA achieving an RFS of 7.556, HMVLM 7.736, and Poutine 7.986. The primary differences among these methods stem from:

### v3

@@ 4.3 Benchmark Models
Swin-Trajectory [ParkSwinTrajectoryTR] is a lightweight, MLP-based model. It uses a Swin Transformer [liu2021swin] to extract image features from three front cameras and a simple MLP to directly predict waypoints. The model is lightweight and achieves a slightly better RFS (7.543) than the baseline.
DiffusionLTF and UniPlan are both Diffusion-based models built on the DiffusionDrive [liao2025diffusiondrive] architecture. Their primary difference lies in the training datasets used: DiffusionLTF utilizes WOD-E2E, CARLA [dosovitskiy2017carla], NAVSIM [Dauner2024NEURIPS], and WOD-Perception [sun2020scalability], whereas UniPlan is trained on WOD-E2E and nuPlan [caesar2021nuplan]. They achieve comparable performance, with RFS scores of 7.717 and 7.779, respectively.
Poutine [pal2025poutinevisionlanguagetrajectorypretraining], HMVLM [wang2025hmvlm], and AutoVLA [zhou2025autovla] are all MLLM-based models that use Qwen2.5 as their backbone. They share a similar problem formulation, taking camera images and ego states as input modalities and outputting future waypoints as text. Additionally, all three models utilize Chain-of-Thought (CoT) reasoning before generating a trajectory. Despite these similarities, their results show a significant performance gap, with AutoVLA achieving an RFS of 7.556, HMVLM 7.736, and Poutine 7.986. The primary differences among these methods stem from: